An investment company pays 4​% compounded semiannually. You want to have $19,000 in the future. How much should you deposit now to have that amount 5 years from​ now?

P​ = ​$ ?

​(Do not round until the final answer. Then round to two decimal places as​ needed.)
Answer #1

We use the formula:
A=P(1+r/2)^2n
where
A=future value
P=present value
r=rate of interest
n=time period.

19000=P*(1+0.04/2)^(2*5)

P=19000/(1+0.04/2)^(2*5)

=19000*0.8203483

=$15586.62(Approx).